# Littlewood, Paley and Almost-Orthogonality: a theory well ahead of its time

### v1

@@ 7. Almost-orthogonality principles beyond Littlewood–Paley
Theorem 7.1 (Cotlar–Stein, see [SteinBig]).
@@
The proof is by a clever but elementary combinatorial argument. The hypothesis is that we have $\sup_{j}\|T_{j}\|<\infty$, plus some almost-orthogonality between the different $T_{j}$. Generally one has an operator $T$ at hand whose boundedness one wants to establish; the art is in finding a natural decomposition $T=\sum_{j}T_{j}$ (if one exists) to which one may apply the Cotlar–Stein lemma. In the case of the Hilbert transform (5) we can take $T_{j}f(x)=\frac{1}{\pi}\int_{|y|\sim 2^{j}}\frac{f(x-y)}{y}{\rm d}y$. Even when a spectral decomposition of $T$ is not readily available, such real-variables decompositions often work provided that $T$ satisfies some cancellation and smoothness properties. Boundedness on $L^{2}$ of various classes of pseudodifferential operators of Calderón–Vaillancourt-type was amongst the first successes of the Cotlar–Stein lemma, see [SteinBig].
The hypothesis can be verified in concrete situations when $H=L^{2}$ as follows. By Schur’s lemma it suffices to show that, with $\|T\|_{p-p}$ denoting the $L^{p}$ operator norm,
@@
for all $j,k\in\mathbb{Z}$, and this may be directly verified if $T$ has appropriate cancellation and its integral kernel has some smoothness. This works in the case of classical singular integrals when, say, $T1=T^{\ast}1=0$. But even when the integral kernel of $T$ is only a measure, and so has no classical smoothness, the paradigm can still be made to work. The idea is that instead of estimating $\|S\|$ directly by $\|S\|_{1-1}$ (where $S=T_{j}T_{k}^{\ast}$), we instead estimate $\|S\|^{2}=\|S^{\ast}S\|$ by $\|S^{\ast}S\|_{1-1}$ (or by $\|SS^{\ast}\|_{1-1}$). This has more chance of working because the effect of the composition of $S^{\ast}$ with $S$ is typically to spread out and indeed smooth out the integral kernel, as in the case of repeated convolution. The more repetitions of this strategy one carries out, the more likely one is to succeed! For an early manifestation of this idea see the work of Christ, [Christ].
@@
Following a suggestion of Michael Cowling, the Cotlar–Stein lemma can be used in conjunction with Littlewood–Paley theory to give further $L^{p}$-almost-orthogonality principles which are applicable in non-commutative and non-translation-invariant settings, see for example [CarbCotlar]. Moreover, as observed by Christ, such principles may be further adapted to give almost-orthogonality principles for maximal functions, the roots of which lie in the work of Nagel, Stein and Wainger [NSW] on differentiation in lacunary directions. See [MR837527] and [CarbConvex]; see also [MR1809116].
8. Littlewood–Paley theory in Geometric Measure Theory and PDE
@@
Littlewood–Paley theory and quadratic expressions such as the area function were, in their nascence, intimately connected to complex function theory and to boundary value problems for harmonic functions on the upper half-plane. It it therefore unsurprising that they continue to play a key role in the more modern and refined study of boundary value problems for elliptic and parabolic PDE, including those with rough coefficients, and on domains with rough boundaries where the boundary data is in $L^{p}$. The analogue of the Hilbert transform in the classical setting becomes a singular integral – a Cauchy-type integral – associated to a rough (hyper-)surface, [^10] It is to be noted that this is a quite different generalisation from $H_{\Gamma}$ discussed above. and the study of such involves the development of singular integral theory going beyond the treatment of David, Journé and Semmes. Calderón established the boundedness of the Cauchy integral along curves with small Lipschitz constant, and Coifman, McIntosh and Meyer did so for arbitrary Lipschitz curves. David, Jones and Semmes pioneered the study of still rougher curves and surfaces. The geometric measure theory of such surfaces begins to play a critical role: boundedness of such a singular integral operator turns out to be closely related to uniform rectifiability of the underlying surface. Quadratic expressions, along with Carleson measures, feature throughout the development, including in the celebrated solution of the Kato square root problem by Auscher, Hofmann, Lacey, McIntosh and Tchamitchian, pushing the theory of singular integrals to its limits, see for example [Hofmann]; and also Tolsa’s solution of the Painlevé problem [^11] The Painlevé problem was one of the initial motivations behind the study of the Cauchy integral on Lipschitz graphs and uniformly rectifiable sets., see [Tolsa]. See also [Verdera], Sections 7 and 8.
@@
Quadratic expressions involving Peter Jones’ $\beta$-numbers and their relatives are central in the study of uniform rectifiability. They also give an analytic perspective on structure in large data sets – in particular measuring clustering of data around rectifiable sets. See [DavidSemmes] (Section I.1.3) which directly addresses the analogies between Littlewood–Paley techniques in analysis and in geometry. There is also a very nice survey [Mattila] by Mattila on rectifiability, see especially Chapter 5 on uniform rectifiability.
The survey article of Dindoš and Pipher [DindosPipher] is a good reference for the modern theory for elliptic PDE with rough coefficients in Lipschitz domains in which quadratic expressions play a crucial role. See also [MR3952696] for connections between the PDE perspective and $\beta$-numbers in higher-codimensional settings.
@@
There is now a vast literature on this area which we can make no attempt to summarise. This section is thus perhaps even more impressionistic than the others. Amongst the contributors in the current era are Azzam, Dąbrowski, David, Dindoš, Feneuil, Hofmann, Kenig, Li, Mattila, Mayboroda, Martell, Mourgoglou, Pipher, Tolsa, Toro, Verdera, Villa and their collaborators.
@@ 9.1. Equally-spaced intervals
It is easily seen that any inequality of the form $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ for $p<2$ with $N$ operators $Q_{j}$ fails by a factor $\sim N^{\frac{1}{p}-\frac{1}{2}}$, (and hence that there is no inequality of the form $\|f\|_{p}\lesssim\|\mathcal{Q}f\|_{p}$ for $p>2$), but Carleson showed that $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ does indeed hold for $p\geq 2$. Later, Córdoba and separately Rubio de Francia gave alternative arguments for this, with Rubio de Francia’s argument [RdeFequal] being especially enlightening: using nothing more than Parseval’s identity and basic properties of the Fourier transform, he showed that for the smooth version of $\mathcal{Q}$, there is essentially an identity (modulo negligible error terms)
@@
where $\Phi$ is a suitably chosen normalised non-negative bump function. Anything that can be said about $\mathcal{Q}$ can be readily derived from this identity. These “equally-spaced" Littlewood-Paley inequalities turn out to be important in problems concerning $L^{p}$-Fourier inversion in $\mathbb{R}^{n}$, see below.
@@ 9.2. Arbitrary intervals – the Rubio de Francia Littlewood–Paley inequality
Shortly thereafter, Rubio de Francia [RdeFarb] further surprised the harmonic analysis community when he proved that one may replace the congruent intervals [j1/2,j+1/2)[j-1/2,j+1/2) by any disjoint collection of intervals whatsoever, and still conclude that $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ for $2\leq p<\infty$. Moreover, in this case, the simple identity above is replaced by a pointwise inequality
@@
where #\# represents the Fefferman–Stein sharp function which appears in the definition of the space BMO and which satisfies $\|F\|_{p}\lesssim\|F^{\#}\|_{p}$ for $1<p<\infty$. (See [SteinBig].) Remarkably, Rubio de Francia’s pointwise inequality is proved by establishing kernel estimates for $f\mapsto\mathcal{Q}(f)$ as a (non-symmetric) vector-valued Calderón–Zygmund operator; this is in spite of the fact that an arbitrary collection of disjoint intervals contains no vestige of dyadic block or dilation structure.
Rubio de Francia’s inequality was later extended to arbitrary families of disjoint axis-parallel rectangles in $\mathbb{R}^{n}$. See [Journe], [Lacey] and also [Roncaletal].
@@ 9.3. Angular decompositions – reverse inequalities and curvature
Notice that the parabolic dilation $\xi\mapsto(2^{k}\xi_{1},2^{2k}\xi_{2})$ maps $\Gamma_{j}$ to $\Gamma_{j+k}$, and this suggests that the Littlewood–Paley theory for lacunary angular sectors might be largely accessible within the established theory. Indeed, the Littlewood–Paley inequalities for smoothed versions of the lacunary sectors $\Gamma_{j}$ are as expected, and are a straightforward consequence of the Marcinkiewicz multiplier theorem. If we want to consider quadratic expressions corresponding to the rough $\Gamma_{j}$, we will need some auxiliary vector-valued and maximal inequalities corresponding to differentiation in lacunary directions [NSW].
The theory for the equiangular sectors $\Delta_{j}$, or their smooth variants, where we see an underlying rotational action, is more challenging and interesting. One does not expect to have estimates of the form $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ for $p<2$ for the same reasons that the equally-spaced interval inequalities fail. But when $p>2$, “pure" estimates with bounds indepenedent of $N$ are not available either, due to Kakeya set examples. [^12] More on this below. However, there is a dimension-dependent window of $p$’s in which one might hope for bounds with at most subpolynomial or logarithmic failure in $N$. In two dimensions this window is $2\leq p\leq 4$, and in higher dimensions it is $2\leq p\leq 2n/(n-1)$. Córdoba [Cordoba] established the positive result in the two-dimensional case, while the higher-dimensional case awaits resolution of the Kakeya maximal conjecture, amongst other things, and remains open.
Just when it seems we have said everything that can currently be said, matters take an interesting turn. One of the most powerful drivers for research in harmonic analysis over the last fifty years has been the question of inverting the Fourier transform for functions in $L^{p}(\mathbb{R}^{n})$ (as opposed to the almost-everywhere questions in Sections 1 and 10). The fundamental question here is, with
@@
whether $S_{R}f\to f$ in $L^{p}$-norm for all $f\in L^{p}(\mathbb{R}^{n})$, which, with a little elementary functional analysis, is easily seen to be equivalent to boundedness on $L^{p}(\mathbb{R}^{n})$ of $S_{R}$, or indeed of $S_{1}$. Fefferman [Feffermanball] astonished the world of harmonic analysis when he showed that this boundedness failed when $n\geq 2$, except when $p=2$. And moreover what is responsible for this failure is the rough singularity of the multiplier $\chi_{|\xi|\leq 1}$ for $S_{1}$ at $|\xi|=1$, together with the fact that the normal directions to the sphere fill an open set. This allowed him to construct counterexamples based on Kakeya sets. For this reason, and others, it is natural to decompose the rough disc multiplier $\chi_{\{|\xi|<1\}}$ into smooth (Littlewood–Paley!) dyadic pieces $\phi_{2^{j}}$ for $j\geq 0$, each supported in an annulus of width $\sim 2^{-j}$ and distant $\sim 2^{-j}$ from the unit sphere, and to ask about $L^{p}$ bounds for the corresponding operators $S^{2^{-j}}$. [^13] The question of $L^{p}$ orthogonality between the different $S^{2^{-j}}$ is deliberately not addressed here. Obviously we have $\|S^{2^{-j}}\|_{2-2}\lesssim 1$ uniformly in $j$. Fix $j$, and let $\delta=2^{-j}$. It is also natural to break up each $\phi_{\delta}$ into $\sim\delta^{-(n-1)/2}$ equiangular smooth pieces $\phi_{\nu}$, each supported in a slab of size $\sim\delta\times\delta^{1/2}\times\dots\times\delta^{1/2}$ contained in the $\delta$-neighbourhood of the sphere, each of which has corresponding operator $S_{\nu}$ which is easily seen to satisfy $\|S_{\nu}\|_{p-p}\lesssim 1$ for $1\leq p\leq\infty$. One quickly concludes using interpolation and the triangle inequality that $\|S^{\delta}\|_{p-p}\lesssim\delta^{-(n-1)|\frac{1}{p}-\frac{1}{2}|}$. However, this is far from the whole story, and the use of the triangle inequality turns out to be a very poor choice. Indeed, there is a huge amount of orthogonality between the different $S_{\nu}$ (remember $\delta$ is fixed here) which we have not thus far exploited. This orthogonality finds its formulation, when $n=2$, in the reverse Littlewood–Paley inequality
$\|\sum_{\nu}S_{\nu}f\|_{4}\lesssim\|(\sum_{\nu}|S_{\nu}f|^{2})^{1/2}\|_{4}.$
Fefferman proved this in [FeffermanIsrael]. At first it appears to run counter to the idea that we can have no equally-spaced forward Littlewood–Paley inequalities for any $p<2$, but on closer inspection it does not violate that principle. Fefferman’s argument, which is very simple, exploits the fact that $4$ is an even integer to multiply out the integral implicit on the left hand side to turn it into an $L^{2}$ expression for which we can use Plancherel’s theorem, and then, crucially, uses the curvature of the circle to observe that the algebraic sums
@@
have bounded overlap. This, together with arguments for the two-dimensional Kakeya maximal operator and the forward Littlewood–Paley inequalities for equally-spaced intervals when $p\geq 2$, turns out to be enough to establish the sharp inequality
@@
Let us emphasise that all this is specific to two dimensions; all of the analogous higher-dimensional questions remain open.
We seem to have lost ourselves down a very specialised and esoteric rabbit hole. But I now wish to suggest that this is far from the case. Three of the most important open questions in euclidean harmonic analysis are: (i) $L^{p}$ bounds for the Kakeya maximal operator (the corresponding dimensional consequence of these when $n=3$ has recently been established in spectacular work of Wang and Zahl [WangZahl]), (ii) $L^{p}$ bounds for the Bochner–Riesz operators $S^{\delta}$, and (iii) $L^{p}-L^{q}$ bounds for the Fourier restriction and extension operators for the sphere. These are known to be intimately related, and all of them are resolved when $n=2$, but all are open in all higher dimensions. Moreover, it is known [carberyRLP] that all three would follow from the higher-dimensional analogue
@@
of Fefferman’s reverse Littlewood–Paley inequality. This inequality (11), if true, would therefore be very deep. Note that when $n\geq 3$, the index 2n/(n1)2n/(n-1) is no longer an even integer.
@@
holds, perhaps up to subpolynomial factors in $\delta^{-1}$. This had resisted all attempts (see [MSS]) for forty years, and has only recently been resolved in a tour de force by Guth, Wang and Zhang in [GuthWangZhang]. The techniques apply not only to the $L^{4}$-boundedness of the operator $\sum_{\nu}S_{\nu}$, but also to local smoothing estimates for dispersive PDE.
@@
(up to sub-polynomial factors in $\delta^{-1}$) for $p\geq 2$. The right-hand side is larger than $\|(\sum_{\nu}|S_{\nu}f|^{2})^{1/2}\|_{p}$ , so there is more scope for establishing positive results, and the range of $p$ for which such might hold is larger than in the classical Littlewood–Paley setting. Indeed, when $n=2$, the critical $p$ turns out now to be the $L^{2}$-extension index $6$ instead of $4$. Wolff [Wolff] established sharp estimates of this sort for large $p$ (with the correct power of $\delta$), and that they could be applied to get sharp $L^{p}$ estimates for radial Fourier multiplier operators for large $p$. See also [MR2078706] and [MR2664568]. Since then, such estimates, now known as decoupling estimates, have been systematically studied and developed, initially by Bourgain and Demeter [BourgainDemeter] (see also [Demeter]), and have had a huge range of applications, not least of which is the resolution of the Main Conjecture in Vinogradov’s Mean Value Theorem, see [BDG].

### v2

@@ 7. Almost-orthogonality principles beyond Littlewood–Paley
Theorem 7.1 (Cotlar–Stein, see [71]).
@@
The proof is by a clever but elementary combinatorial argument. The hypothesis is that we have $\sup_{j}\|T_{j}\|<\infty$, plus some almost-orthogonality between the different $T_{j}$. Generally one has an operator $T$ at hand whose boundedness one wants to establish; the art is in finding a natural decomposition $T=\sum_{j}T_{j}$ (if one exists) to which one may apply the Cotlar–Stein lemma. In the case of the Hilbert transform (5), for example, it is natural to take $T_{j}f(x)=\frac{1}{\pi}\int_{|y|\sim 2^{j}}\frac{f(x-y)}{y}{\rm d}y$ (of course we do not need to do this, because we already know the Hilbert transform is bounded). Even when a spectral decomposition of $T$ is not readily available, such real-variables decompositions often work provided that $T$ satisfies some cancellation and smoothness properties. Boundedness on $L^{2}$ of various classes of pseudodifferential operators of Calderón–Vaillancourt-type was amongst the first successes of the Cotlar–Stein lemma, see [71].
The hypothesis can be verified in certain concrete situations when $H=L^{2}$ as follows. By Schur’s lemma it suffices to show that, with $\|T\|_{p-p}$ denoting the $L^{p}$ operator norm,
@@
for all $j,k\in\mathbb{Z}$, and this may be directly verified if $T$ has appropriate cancellation and its integral kernel has some smoothness. This works in the case of classical non-convolution singular integral operators $T$, when, say, $T1=T^{\ast}1=0$. But even when the integral kernel of $T$ is only a measure, and so has no classical smoothness, the technique can still be made to work. Let $S=T_{j}T_{k}^{\ast}$ or $T_{j}^{\ast}T_{k}$. The idea is that instead of estimating $\|S\|$ directly by $\|S\|_{1-1}$, we instead estimate $\|S\|^{2}=\|S^{\ast}S\|$ by $\|S^{\ast}S\|_{1-1}$ (or by $\|SS^{\ast}\|_{1-1}$). This has more chance of working because the effect of the composition of $S^{\ast}$ with $S$ is typically to spread out and smooth out the integral kernel, as in the case of repeated convolution. The more repetitions of this strategy one carries out, the more likely one is to succeed! For an early manifestation of this idea see the work of Christ, [21].
@@
Following a suggestion of Michael Cowling, the Cotlar–Stein lemma can be used in conjunction with Littlewood–Paley theory to give further $L^{p}$-almost-orthogonality principles which are applicable in non-commutative and non-translation-invariant settings, see for example [11]. Moreover, as observed by Christ, such principles may be further adapted to give almost-orthogonality principles for maximal functions, the roots of which lie in the work of Nagel, Stein and Wainger [60] on differentiation in lacunary directions. See [34] and [15]; see also [20].
8. Littlewood–Paley theory in PDE and Geometric Measure Theory
Littlewood–Paley theory and quadratic expressions – especially those such as Lusin’s area function – were, in their nascence, intimately connected to complex function theory and to boundary value problems for harmonic functions on the upper half-plane. It it therefore unsurprising that they continue to play a key role in the more modern and refined study of the solvability of boundary value problems for elliptic and parabolic PDE, including those with rough coefficients, and on domains with rough boundaries where the boundary data is in $L^{p}$. The quadratic expressions which occur are, naturally, often more closely related to the Lusin area function than to the more Fourier-analytic square functions that we have focused on in this article. However, Littlewood–Paley expressions associated to the diffusion semigroup generated by the uniformly elliptic operator ${\rm div}(A\nabla)$ are also important. The survey article of Dindoš and Pipher [33] is a good reference for the modern theory for uniformly elliptic PDE with rough coefficients in Lipschitz domains in which quadratic expressions feature strongly.
@@ 8. Littlewood–Paley theory in PDE and Geometric Measure Theory
The analogue of the Hilbert transform in the classical setting becomes a singular integral – a Cauchy-type integral – associated to a rough (hyper-)surface, [^11] It is to be noted that this is a quite different generalisation from $H_{\Gamma}$ discussed above. and the study of such involves the development of singular integral theory going beyond the treatment of David, Journé and Semmes. Calderón established the boundedness of the Cauchy integral along curves with small Lipschitz constant, and Coifman, McIntosh and Meyer did so for arbitrary Lipschitz curves. David, Jones and Semmes pioneered the study of still rougher curves and surfaces. The geometric measure theory of such surfaces begins to play a critical role: boundedness of such a singular integral operator turns out to be closely related to uniform rectifiability of the underlying surface. Quadratic expressions, along with Carleson measures, feature throughout the development, including in the celebrated solution of the Kato square root problem by Auscher, Hofmann, Lacey, McIntosh and Tchamitchian, pushing the theory of singular integrals to its limits, see for example [43]; and also Tolsa’s solution of the Painlevé problem [^12] The Painlevé problem was one of the initial motivations behind the study of the Cauchy integral on Lipschitz graphs and uniformly rectifiable sets., see [72]. See also [73], Sections 7 and 8.
@@
Quadratic expressions involving Peter Jones’ $\beta$-numbers and their relatives are central in the study of uniform rectifiability. They also give an analytic perspective on structure in large data sets – in particular measuring clustering of data around rectifiable sets. See [30] (Section I.1.3) which directly addresses the analogies between Littlewood–Paley techniques in analysis and in geometry. There is also a very nice survey [56] by Mattila on rectifiability, see especially Chapter 5 on uniform rectifiability. See [28] for connections between the PDE perspective and $\beta$-numbers in higher-codimensional settings.
@@
There is now a vast literature on this area which we can make no attempt to summarise; partly for this reason this section is thus perhaps even more impressionistic than the others. Amongst the contributors in the current era are Azzam, Dąbrowski, David, Dindoš, Feneuil, Hofmann, Kenig, Li, Martell, Mattila, Mayboroda, Mourgoglou, Pipher, Tolsa, Toro, Verdera, Villa and their collaborators.
@@ 9.1. Equally-spaced intervals
It is easily seen that any inequality of the form $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ for $p<2$ with $N$ operators $Q_{j}$ fails by a factor $\sim N^{\frac{1}{p}-\frac{1}{2}}$, (and hence that there is no inequality of the form $\|f\|_{p}\lesssim\|\mathcal{Q}f\|_{p}$ for $p>2$), but Carleson showed that $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ does indeed hold for $p\geq 2$. Later, Córdoba and separately Rubio de Francia gave alternative arguments for this, with Rubio de Francia’s argument [64] being especially informative: using nothing more than Parseval’s identity and basic properties of the Fourier transform, he showed that for the smooth version of $\mathcal{Q}$, there is essentially an identity (modulo negligible error terms)
@@
where $\Phi$ is a suitably chosen normalised non-negative bump function. Anything that can be said about $\mathcal{Q}$ can be readily derived from this relation. Perhaps surprisingly, these “equally-spaced" Littlewood-Paley inequalities turn out to be important in problems concerning $L^{p}$-Fourier inversion in $\mathbb{R}^{n}$, see below.
@@ 9.2. Arbitrary intervals – the Rubio de Francia Littlewood–Paley inequality
Shortly thereafter, Rubio de Francia [65] further surprised the harmonic analysis community when he proved that one may replace the congruent intervals [j1/2,j+1/2)[j-1/2,j+1/2) considered above by any disjoint collection of intervals whatsoever, and still conclude that $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ for $2\leq p<\infty$. Moreover, in this case, his simple identity is replaced by a pointwise inequality
@@
where #\# represents the Fefferman–Stein sharp function which appears in the definition of the space BMO and which satisfies $\|F\|_{p}\lesssim\|F^{\#}\|_{p}$ for $1<p<\infty$. (See [71].) Remarkably, Rubio de Francia’s pointwise inequality is proved by establishing kernel estimates for $f\mapsto\mathcal{Q}(f)$ as a (non-symmetric) vector-valued Calderón–Zygmund operator; this is in spite of the fact that an arbitrary collection of disjoint intervals contains no vestiges of dyadic block or dilation structure.
Rubio de Francia’s inequality was later extended to arbitrary families of disjoint axis-parallel rectangles in $\mathbb{R}^{n}$. See [45], [49] and also [32].
@@ 9.3. Angular decompositions – reverse inequalities and curvature
Notice that the parabolic dilation $\xi\mapsto(2^{k}\xi_{1},2^{2k}\xi_{2})$ maps $\Gamma_{j}$ to $\Gamma_{j+k}$, and this suggests that the Littlewood–Paley theory for lacunary angular sectors might be largely accessible within the established theory. Indeed, the Littlewood–Paley inequalities for smoothed versions of the lacunary sectors $\Gamma_{j}$ are as expected, and are a straightforward consequence of the Marcinkiewicz multiplier theorem. If we want to consider quadratic expressions corresponding to the rough $\Gamma_{j}$, we will need some auxiliary vector-valued and maximal inequalities corresponding to differentiation in lacunary directions [60].
The theory for the equiangular sectors $\Delta_{j}$, or their smooth variants, where we see an underlying rotational action, is more challenging and interesting. One does not expect to have estimates of the form $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ for $p<2$ for the same reasons that the inequalities for equally-spaced intervals fail. But when $p>2$, “pure" estimates with bounds independent of $N$ are not available either, due to Kakeya set examples. [^13] More on this below. However, there is a dimension-dependent window of $p$’s with $p>2$ in which one might hope for bounds of the form $\|\mathcal{Q}f\|_{p}\lesssim\|f\|_{p}$ with at most subpolynomial or logarithmic failure in $N$. In two dimensions this window is $2\leq p\leq 4$, and in higher dimensions it is $2\leq p\leq 2n/(n-1)$. Córdoba [23] established the positive result in the two-dimensional case, while the higher-dimensional case awaits resolution of the Kakeya maximal conjecture, amongst other things, and remains open.
Just when it seems we have said everything that can currently be said, matters take an interesting turn. One of the most powerful drivers for research in harmonic analysis over the last fifty years has been the question of inverting the Fourier transform for functions in $L^{p}(\mathbb{R}^{n})$ (as opposed to the almost-everywhere questions of Sections 1 and 10). The fundamental question here is, with
@@
whether $S_{R}f\to f$ in $L^{p}$-norm for all $f\in L^{p}(\mathbb{R}^{n})$, (say for $p\leq 2$), which, with a little elementary functional analysis, is easily seen to be equivalent to boundedness on $L^{p}(\mathbb{R}^{n})$ of the operator $S_{R}$, or indeed of $S_{1}$. Fefferman [35] astounded the world of harmonic analysis when he showed that this boundedness failed when $n\geq 2$, except when $p=2$. And moreover what is responsible for this failure is the rough singularity of the multiplier $\chi_{|\xi|\leq 1}$ for $S_{1}$ at $|\xi|=1$, together with the fact that the normal directions to the sphere fill an open set. This allowed him to construct counterexamples based on Kakeya sets. For this reason, and others, it is natural to decompose the rough disc multiplier $\chi_{\{|\xi|<1\}}$ into smooth (Littlewood–Paley!) dyadic pieces $\phi_{2^{j}}$ for $j\geq 0$, each supported in an annulus of width $\sim 2^{-j}$ and distant $\sim 2^{-j}$ from the unit sphere, and to ask about $L^{p}$ bounds for the corresponding operators $S^{2^{-j}}$. [^14] The question of $L^{p}$ orthogonality between the different $S^{2^{-j}}$ is an interesting matter which we will pursue elsewhere. Obviously we have $\|S^{2^{-j}}\|_{2-2}\lesssim 1$ uniformly in $j$. Fix $j$, and let $\delta=2^{-j}$. It is also natural to break up each $\phi_{\delta}$ into $\sim\delta^{-(n-1)/2}$ equiangular smooth pieces $\phi_{\nu}$, each supported in a slab of size $\sim\delta\times\delta^{1/2}\times\dots\times\delta^{1/2}$ contained in the $\delta$-neighbourhood of the sphere, each of which has corresponding operator $S_{\nu}$ which is easily seen to satisfy $\|S_{\nu}\|_{p-p}\lesssim 1$ for $1\leq p\leq\infty$. One quickly concludes using interpolation and the triangle inequality that $\|S^{\delta}\|_{p-p}\lesssim\delta^{-(n-1)|\frac{1}{p}-\frac{1}{2}|}$. However, the use of the triangle inequality here turns out to be a very poor choice, and we can do much better. Indeed, there is a huge amount of orthogonality between the different $S_{\nu}$ (remember $\delta$ is fixed here) which we have not thus far exploited. This orthogonality finds its formulation, when $n=2$, in the reverse Littlewood–Paley inequality
$\|\sum_{\nu}S_{\nu}f\|_{4}\lesssim\|(\sum_{\nu}|S_{\nu}f|^{2})^{1/2}\|_{4},$
which Fefferman proved in [36]. At first this appears to run counter to the observation that we can have no equally-spaced forward Littlewood–Paley inequalities for any $p<2$, but on closer inspection it does not violate that principle. Fefferman’s argument, which is very simple, exploits the fact that $4$ is an even integer to multiply out the integral implicit on the left hand side to turn it into an $L^{2}$ expression for which we can use Plancherel’s theorem, and then, crucially, uses the curvature of the circle to observe that the algebraic sums
@@
have bounded overlap. This, together with arguments for the two-dimensional Kakeya maximal operator and the forward Littlewood–Paley inequalities for equally-spaced intervals when $p\geq 2$ which we discussed above, turns out to be enough to establish the sharp inequality
@@
Let us emphasise that all this is specific to two dimensions, see [22]; the analogous higher-dimensional questions remain open.
Indeed, three of the most important open questions in euclidean harmonic analysis on $\mathbb{R}^{n}$ for $n\geq 3$ are: (i) $L^{p}$ bounds for the (Bochner–Riesz) operators $S^{\delta}$, (ii) $L^{p}$ bounds for the Kakeya maximal operator (although the corresponding assertion that Kakeya sets have full dimension when $n=3$ has recently been established in spectacular work of Wang and Zahl [74]), and (iii) $L^{p}-L^{q}$ bounds for the Fourier restriction and extension operators for the sphere. These three questions are known to be intimately related, and all of them are resolved when $n=2$, but all are open in all higher dimensions. Moreover, it is known [18] that all three would follow from the higher-dimensional analogue
@@
of Fefferman’s reverse Littlewood–Paley inequality. This inequality (11), if true, would therefore be profound. Note that when $n\geq 3$, the index 2n/(n1)2n/(n-1) is no longer an even integer.
@@
holds, perhaps up to subpolynomial factors in $\delta^{-1}$. Despite the fact that $4$ is an even integer, this problem had resisted all attempts (see [59]) for forty years, and has only recently been resolved in a tour de force by Guth, Wang and Zhang in [40]. The techniques apply not only to the $L^{4}$-boundedness of the operator $\sum_{\nu}S_{\nu}$, but also to local smoothing estimates for dispersive PDE.
@@
(up to sub-polynomial factors in $\delta^{-1}$) for $p\geq 2$. The right-hand side is larger than $\|(\sum_{\nu}|S_{\nu}f|^{2})^{1/2}\|_{p}$ , so there is more scope for establishing positive results, and the range of $p$ for which such might hold is larger than in the classical Littlewood–Paley setting. Indeed, when $n=2$, the critical $p$ turns out now to be the $L^{2}$-extension index $6$ instead of $4$. Wolff [75] established sharp estimates of this sort for large $p$ (with the correct power of $\delta$), and that they could be applied to get sharp $L^{p}$ estimates for radial Fourier multiplier operators for large $p$. See also [3] and [38]. Since then, such estimates, now known as decoupling estimates, have been systematically studied and developed, initially by Bourgain and Demeter [8] (see also [31]), and have had a huge range of applications, not least of which is the resolution of the Main Conjecture in Vinogradov’s Mean Value Theorem, see [7].